Two spheres of mass 1.0 and 4.0kg are joined by a light rod so that their centre are 0.20m apart. Calculate The position of centre of gravity


1
Expert's answer
2021-08-16T08:36:08-0400

Let the 1.0 kg sphere is placed at the origin. Then

"x_1=0,\quad x_2=0.20\:\rm m"

The position of centre of gravity

"x_c=\frac{x_1m_1+x_2m_2}{m_1+m_2}\\
\small=\frac{0*1.0+0.20*4.0}{1.0+4.0}=0.16\:\rm m\quad\small from\: 1.0\: kg\; sphere"